OpenStudy (anonymous):

rewrite x^-1/3 with positive exponents

OpenStudy (anonymous):

\[x^{\frac{-1}{3}} =\frac {1}{x^{\frac{1}{3}}}\]

OpenStudy (anonymous):

that is what I thought but was not sure.

OpenStudy (anonymous):

\[\sqrt{6t}\]\[\sqrt{6t}\]. multiply then simplify by factoring. Assume all variables under radicals represent non negative numbers.

OpenStudy (anonymous):

\[\sqrt{6t} \times \sqrt{6t} = \sqrt{36t^2}=6t\]

OpenStudy (anonymous):

okay so I just multiply everything in the problem. that makes sense. so if I had a problem that said find the function values.. g(89)= \[\sqrt[4]{t-8}\] would my answer be 9?

OpenStudy (anonymous):

pls wait

OpenStudy (anonymous):

it would be 3

OpenStudy (anonymous):

since the order of the given radical is 4

OpenStudy (anonymous):

okay so if g(-1) it would be 3. g(6569) i got 81 or do I take it down to 3 as well. g(9) I got 1. man I am so lost.